# Preasymptotic error estimates of higher-order EEM for the time-harmonic Maxwell equations with large wave number

### v1

@@ 5. Numerical example
In this section, we simulate the following three-dimensional time-harmonic Maxwell problem to verify our theoretical results:
$\begin{cases}\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}\bm{E}-\kappa^{2}\bm{E}=\bm{f},\quad&{\rm in}\quad\Omega:={(0,1)}^{3},\\ \operatorname{\bm{\mathrm{curl}}}\bm{E}\times\bm{\nu}-\bm{\mathrm{i}}\kappa\bm{E}_{T}=\bm{g},\quad&{\rm on}\quad\Gamma:=\partial\Omega,\\ \end{cases}$
and $\bm{f}$ and $\bm{g}$ are so chosen that the exact solution is $\bm{E}=\big{(}\sin(\kappa y)J_{0}(\kappa r),\cos(\kappa z)J_{0}(\kappa r),\bm{\mathrm{i}}\kappa J_{0}(\kappa r)\big{)}^{T}$, where $r=(x^{2}+y^{2}+z^{2})^{\frac{1}{2}}$ and $J_{0}$ is the Bessel function of the first kind. Given a positive integer $M>0$, the computational domain is discretized first by dividing it into small cubes with side-length $h_{0}:=1/M$, then each small cube is further divided in to six tetrahedrons (see Figure LABEL:fig:mesh). A simple calculation yields that the number of degrees of freedom of the linear system (2.5) is $\text{DOF}:=M(p+1)(3M^{2}p^{2}+3M^{2}p+M^{2}+6Mp+3M+3).$ Similar to [43, Section 10], we denote by $N_{\lambda}$ the number of degrees of freedom per wavelength, which can be calculated by
@@
then we follow [36, Theorem 3.2] to derive (3.14) and (3.15).
@@
${\bm{w}}=\bm{\Upsilon}^{0}+\nabla\psi,$
where $\bm{\Upsilon}^{0}$ satisfies ${\operatorname{\bm{\mathrm{curl}}}}\bm{\Upsilon}^{0}=\operatorname{\bm{\mathrm{curl}}}{\bm{w}}$; By using the fact that ${\bm{w}}\in{\bm{H}}(\operatorname{{\mathrm{div}}}^{0};\Omega)$ and $\bm{\Upsilon}^{0}\in{\bm{H}}_{0}(\operatorname{{\mathrm{div}}}^{0};\Omega)$, $\psi$ can be chosen to satisfy
$\left\{\begin{array}[]{l}\Delta~{}\psi=0,\ \ \ \ \ {\rm in}\ \ \Omega,\\ \displaystyle\frac{\partial\psi}{\partial\bm{\nu}}={\bm{w}}\cdot\bm{\nu},\ \ \ \ \text{ on }\ \ \Gamma,\\ \displaystyle\int_{\Omega}\psi\,{\rm d}x=0.\end{array}\right.$
According to the orthogonality of the Helmholtz decomposition, it holds that
$\left\|\bm{\Upsilon}^{0}\right\|\leq\|{\bm{w}}\|\quad\mbox{and}\quad\|\nabla\psi\|\leq\|{\bm{w}}\|.$
@@
$\left\|\bm{\Upsilon}^{0}\right\|_{{\bm{H}}^{1}(\Omega)}\lesssim\|\operatorname{\bm{\mathrm{curl}}}{\bm{w}}\|.$ (A.2)
@@
$\|\psi\|_{1}\lesssim\|\nabla\psi\|\lesssim\|{\bm{w}}\|_{{\bm{L}}^{2}(\Omega)}.$ (A.3)
@@
where $\nabla_{\Gamma}\psi$ is the tangential gradient of $\psi$, i.e., $\nabla_{\Gamma}\psi=(\bm{\nu}\times\nabla\psi)\times\bm{\nu}$. Recall the formula $\operatorname{{\mathrm{div}}}_{\Gamma}(\bm{v}\times\bm{\nu})=\bm{\nu}\cdot\operatorname{\bm{\mathrm{curl}}}\bm{v}$, we have
@@
$\left\|\bm{\mathrm{i}}\kappa\lambda\bm{\Upsilon}^{0}_{T}\right\|_{\frac{1}{2},\Gamma}\lesssim\kappa\|\bm{\Upsilon}^{0}\|_{{\bm{H}}^{1}(\Omega)}\lesssim\kappa\|\operatorname{\bm{\mathrm{curl}}}{\bm{w}}\|.$ (A.5)
@@
Hence, applying the elliptic regularity theorem for the Laplace-Beltrami operator on smooth surfaces, we have
$\begin{array}[]{ll}\kappa\|\psi\|_{\frac{3}{2},\Gamma}&\lesssim\kappa\left\|\operatorname{{\mathrm{div}}}_{\Gamma}\nabla_{\Gamma}\psi\right\|_{-\frac{1}{2},\Gamma}+\kappa\|\psi\|_{\Gamma}\\ &\lesssim\kappa\left\|\operatorname{{\mathrm{div}}}_{\Gamma}\nabla_{\Gamma}\psi\right\|_{-\frac{1}{2},\Gamma}+\kappa\|\psi\|_{1}\\ &\lesssim\|{\bm{f}^{\prime\prime}}\|+\kappa^{2}\|{\bm{w}}\|+\kappa\|\operatorname{\bm{\mathrm{curl}}}{\bm{w}}\|+\|\operatorname{{\mathrm{div}}}_{\Gamma}{\bm{g}^{\prime}}\|_{-\frac{1}{2},\Gamma},\end{array}$
@@
$\left\{\begin{array}[]{rlrl}\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}{\bm{w}}-{\bm{w}}&=({\bm{f}^{\prime}}-C_{\mathcal{T}}\kappa^{2}\mathcal{T}\textsf{P}^{\perp}{\bm{w}})+(\kappa^{2}-1){\bm{w}}&{\rm in}\ \Omega,\\ \operatorname{\bm{\mathrm{curl}}}{\bm{w}}\times\bm{\nu}-\bm{\mathrm{i}}{\bm{w}}_{T}&={\bm{g}^{\prime}}+\bm{\mathrm{i}}(\kappa-1){\bm{w}}_{T}&{\rm on}\ \Gamma.\end{array}\right.$
@@
which, combined with (3.13)–(3.14), proves (3.16) for $m=2$. Similar to the proof of (2.19), (3.16) can be proved by considering the higher-order regularity estimate of problem (A.1) and applying induction for $m$. This completes the proof of the lemma.
@@
Let ${\bm{\Phi}}_{h}:=\tilde{{\bm{u}}}_{h}-{\bm{v}}_{h}\in\bm{V}_{h}$. Similar to Lemma 2.4, we have the following decompositions (see, e.g., [46, Remark 3.46, Lemma 7.6]):
@@
Finally, the proof of Lemma 4.2 follows by combining Lemma B.1B.3 and using the fact that $\kappa h$ is sufficiently small. ∎

### v2

@@ 5. Numerical example
In this section, we simulate the following three-dimensional time-harmonic Maxwell problem to illustrate our theoretical results:
[…]
and $\bm{f}$ and $\bm{g}$ are so chosen that the exact solution is $\bm{E}=\big(\sin(\kappa y)J_{0}(\kappa r),\cos(\kappa z)J_{0}(\kappa r),\bm{\mathrm{i}}\kappa J_{0}(\kappa r)\big)^{T}$, where $r=(x^{2}+y^{2}+z^{2})^{\frac{1}{2}}$ and $J_{0}$ is the Bessel function of the first kind. Given a positive integer $M>0$, the computational domain is discretized first by dividing it into small cubes with side-length $h_{0}:=1/M$, then each small cube is further divided in to six tetrahedrons (see Figure LABEL:fig:mesh). A simple calculation yields that the number of degrees of freedom of the linear system (2.5) is $\text{DOF}:=M(p+1)(3M^{2}p^{2}+3M^{2}p+M^{2}+6Mp+3M+3).$ As in [43, Section 10], we denote by $N_{\lambda}$ the number of degrees of freedom per wavelength, which can be calculated by
@@
then we follow the proof of [36, Theorem 3.2] to derive (3.14) and (3.15).
@@
${\bm{w}}=\bm{\Upsilon}^{0}+\nabla\psi.$
It is esay to verify that ${\operatorname{\bm{\mathrm{curl}}}}\bm{\Upsilon}^{0}=\operatorname{\bm{\mathrm{curl}}}{\bm{w}}$, $\left\|\bm{\Upsilon}^{0}\right\|\leq\|{\bm{w}}\|$, $\|\nabla\psi\|\leq\|{\bm{w}}\|$, and $\psi$ can be chosen according to the following elliptic equation in weak sense:
$\Delta~\psi=0\ \ \text{in}\ \Omega,\quad\frac{\partial\psi}{\partial\bm{\nu}}={\bm{w}}\cdot\bm{\nu}\ \ \text{on}\ \ \Gamma,\quad\int_{\Omega}\psi\,{\rm d}x=0.$
@@
$\left\|\bm{\Upsilon}^{0}\right\|_{1}\lesssim\|\operatorname{\bm{\mathrm{curl}}}{\bm{w}}\|.$ (A.2)
@@
$\|\psi\|_{1}\lesssim\|\nabla\psi\|\lesssim\|{\bm{w}}\|.$ (A.3)
@@
where $\nabla_{\Gamma}\psi$ is the tangential gradient of $\psi$, i.e., $\nabla_{\Gamma}\psi=(\bm{\nu}\times\nabla\psi)\times\bm{\nu}{=(\nabla\psi)_{T}}$ (see e.g.[47]). Recall the formula $\operatorname{{\mathrm{div}}}_{\Gamma}(\bm{v}\times\bm{\nu})=\bm{\nu}\cdot\operatorname{\bm{\mathrm{curl}}}\bm{v}$ (see e.g. [47]), we have
@@
$\left\|\bm{\mathrm{i}}\kappa\lambda\bm{\Upsilon}^{0}_{T}\right\|_{\frac{1}{2},\Gamma}\lesssim\kappa\|\bm{\Upsilon}^{0}\|_{1}\lesssim\kappa\|\operatorname{\bm{\mathrm{curl}}}{\bm{w}}\|.$ (A.5)
@@
Hence, applying the elliptic regularity theorem for the Laplace-Beltrami operator on smooth surfaces (see e.g. [19, Remark 1] or [20, §1.3.b]), we have
$\kappa\|\psi\|_{\frac{3}{2},\Gamma}$ $\lesssim\kappa\left\|\operatorname{{\mathrm{div}}}_{\Gamma}\nabla_{\Gamma}\psi\right\|_{-\frac{1}{2},\Gamma}+\kappa\|\psi\|_{\Gamma}$
$\lesssim\kappa\left\|\operatorname{{\mathrm{div}}}_{\Gamma}\nabla_{\Gamma}\psi\right\|_{-\frac{1}{2},\Gamma}+\kappa\|\psi\|_{1}$
$\lesssim\|{\bm{f}^{\prime\prime}}\|+\kappa^{2}\|{\bm{w}}\|+\kappa\|\operatorname{\bm{\mathrm{curl}}}{\bm{w}}\|+\|\operatorname{{\mathrm{div}}}_{\Gamma}{\bm{g}^{\prime}}\|_{-\frac{1}{2},\Gamma},$
@@
$\left\{\begin{array}[]{rlrl}\operatorname{\bm{\mathrm{curl}}}\operatorname{\bm{\mathrm{curl}}}{\bm{w}}-{\bm{w}}&=({\bm{f}^{\prime}}-C_{\mathcal{T}}\kappa^{2}\mathcal{T}\textsf{P}^{\perp}{\bm{w}})+(\kappa^{2}-1){\bm{w}}\qquad&{\rm in}\ \Omega,\\ \operatorname{\bm{\mathrm{curl}}}{\bm{w}}\times\bm{\nu}-\bm{\mathrm{i}}{\bm{w}}_{T}&={\bm{g}^{\prime}}+\bm{\mathrm{i}}(\kappa-1){\bm{w}}_{T}\qquad&{\rm on}\ \Gamma.\end{array}\right$.
@@
which, combined with (3.13)–(3.14), proves (3.16) for $m=2$. Similar to the proof of (2.19), (3.16) can be proved by considering the higher-order regularity estimate of problem (A.1) and applying induction for $m$. We omit the details and conclude the proof of the lemma.
@@
Let ${\bm{\Phi}}_{h}:=\tilde{{\bm{u}}}_{h}-{\bm{v}}_{h}\in\bm{V}_{h}$. Similar to Lemma 2.4, we have the following decompositions (see, e.g., [47, Remark 3.46, Lemma 7.6]):
@@
Finally, the proof of Lemma 3.4 follows by combining Lemmas B.1B.3 and using the fact that $\kappa h$ is sufficiently small. ∎